OpenStudy (anonymous):

how to solve y= -x+3 and how to get x-intercept and y-intercept

OpenStudy (anonymous):

you dont have to solve it .. it represent line y intercept means x = 0 so : y(0) = 3 (0,3) - y intercept now x intercept means y = 0 so : 0 = -x +3 x = 3 (3,0) - x intercept

OpenStudy (anonymous):

ty nd hb y= 3/4x-2 y-intercept nd x-intercept??